### v2

# Eye of the vortex: bound spectra in tunable horizonless rotational analogs

Abstract
Analog gravity experiments are making remarkable strides in unveiling both the classical and quantum nature of black holes. By harnessing diverse states of matter, contemporary tabletop setups now replicate strong-field phenomena typically confined to the enigmatic regions surrounding black holes. Through these modern gravity simulators, physical processes once considered elusive may finally be brought into experimental reach. In this work, we investigate the bound spectrum of massless scalar excitations propagating within the effective geometry of a rotating acoustic metric. Specifically, we utilize an analog vortex endowed with a tunable parameter that emulates the spacetime of a rotating gravitational background. This model accommodates both the presence of a sonic horizon–characteristic of an acoustic black hole–for non-zero tuning parameters, and its absence when the parameter vanishes, yielding a horizonless, purely rotational vortex flow devoid of radial inflow. We focus on the latter case, where the vortex flow is purely rotational, and compute the spectral properties of the analog system. The resulting bound-state spectrum is found to be qualitatively consistent with that observed in recent experimental realizations of superfluid Helium giant quantum vortices featuring solid or hollow cores. This correspondence suggests that the analog spacetime geometry used here holds significant potential to replicate the phenomenology of cutting-edge laboratory experiments. In doing so, it offers new insight into the intricate landscape of analog black hole spectroscopy and, potentially, the physical topography of bounded, rotating astrophysical environments around black holes.
I Introduction
Gravitational-wave (GW) events from coalescing binaries are currently flooding ground-based detectors with data from compact binary sources. The LIGO-Virgo-KAGRA (LVK) collaboration has successfully detected more than $200$ GW events [1]. These confirmed observations are sourced by mostly black hole (BH) and neutron star (NS) mergers, as well as BH-NS binary systems. Their characteristic imprint on the LVK detectors is divided in three stages; an initial inspiral stage, where the two compact objects meander around a center of mass whilst emitting GWs, thus their orbital distance is shrinking, followed by a merger phase, where the two objects plow into each other violently. The final stage is dubbed the ringdown, where the vibrating remnant tends into equilibrium by radiating away its final fluctuations in a series of damped sinusoids, known as the quasinormal modes (QNMs) [2, 3, 4].
Concurrently, analog gravity simulators have shown great progress in the field of classical and quantum BH phenomenology [5, 6, 7, 8, 9, 10]. Contemporary technological advances in controlling and manipulating such emulators have enabled the experimental realization of acoustic analogs of gravitational BHs [11]. The past decade witnessed a large variety of laboratory experiments testing a diverse set of aspects of BH physics that led to tentative evidence of analog Hawking radiation [12, 13, 14, 15, 16, 17], superradiance [18, 19] and imprints of QNMs in perturbed analogs [20, 21, 22]. Bridging the gap between analog BH experiments and realistic GWs is now in full bloom. The latest experimental and theoretical developments in these areas illustrate that there is an overabundance of possibilities, even at cosmological scales [23, 24, 25].
Even though analog gravity was initially proposed by Unruh [11], in order to understand the quantum nature of spacetime through analog BHs with sonic horizons, beyond which acoustic excitation become supersonic and cannot escape the hole, at the moment the only observed phenomenon from binary mergers is the emission of GWs, as described above, which is a purely classical manifestation of strong-field gravity. The study of a remnant’s QNMs, dubbed as the BH spectroscopy program [26, 27, 28, 29], can serve as a validity test for analog BH spectroscopy [20]. Since the experimental results from analog BH laboratories correctly hint the existence of QNMs, and even quasibound states (QBSs) [30], so the observed analog Hawking radiation may provide, at least qualitatively, legitimate evidence for the quantum nature of gravitation [31, 32].
Most BH analogs that are built with fluids [33, 34, 35, 36, 37, 38], optical lasers [39, 40, 41, 42, 43], Bose-Einstein condensates [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], superconductors [55, 56, 57] and even quantum circuits [58, 59, 60, 61] have one thing in common; an acoustic horizon. Recent experiments also utilize superfluids [62, 63, 64, 65, 66, 67], such as superfluid Helium. In some particular superfluid-Helium cases, the acoustic horizon can even be absent, i.e. there is no radial inflow of surface excitations. Recently, such laboratory contraption has been constructed and dubbed a giant quantum vortex [68], due to the enormous amount of vortex quanta used to build the vortex [69]. It has been shown that such experiments, when perturbed with surface waves, produce both QNMs and QBSs, regardless of the absence of an event horizon [68, 22]. Utilizing such a circulating vortex without radial flow, and calculating its spectrum is, nonetheless, absent from the current literature.
In this work we revisit an effective metric that describes acoustic excitations on a rotating geometry, with a tuning parameter that can turn the radial inflow on and off. This geometry can describe both the vortex flow experiments involving superfluid Helium without radial flow [68, 22], and rotating acoustic BH analogs with radial flow. Here, we focus on the case of a horizonless acoustic analog, where radial flow is absent, in an attempt to qualitatively replicate the results presented in [68], regarding the QBSs. We find both the co-rotating and counter-rotating QBSs of the rotational horizonless vortex. These are qualitatively compared with the results in [68]. Our analysis matches very well the qualitative results for co-rotating and counter-rotating QBSs, thus we believe that the theoretical setup used here can help into further elucidating the perturbative phenomenology of rotational horizonless vortices, which otherwise is experimentally sensitive to nonlinear dispersion and mechanical noise.
II Rotating acoustic black holes
We begin by considering the acoustic BH solution in Minkowski spacetime obtained by Unruh [11]. Then we discuss a choice for the fluid flow velocity in order to obtain a rotating acoustic BH [5, 7] in cylindrical coordinates, which could also be named laboratory coordinates. By switching off the tuning parameter related to the inflow, we obtain the acoustic metric describing a purely rotational fluid flow.
The fundamental equations of motion for an irrotational fluid are given by
×$\nabla\times\mathbf{v}$ $=$ $0,$ (1)
$\partial_{t}\rho+\nabla\cdot(\rho\mathbf{v})$ $=$ $0,$ (2)
$\rho[\partial_{t}\mathbf{v}+(\mathbf{v}\cdot\nabla)\mathbf{v}]+\nabla p$ $=$ $0,$ (3)
where $\mathbf{v}$, $\rho$, and $p$ are the velocity, density, and pressure of the fluid, respectively. We introduce the velocity potential $\Psi$, such that $\mathbf{v}=-\nabla\Psi$, and assume the fluid as barotropic, which means that $\rho=\rho(p)$. Then, by linearizing these equations of motion around some background $(\rho_{0},p_{0},\Psi_{0})$, namely,
$\rho$ $=$ $\rho_{0}+\epsilon\rho_{1},$ (4)
$p$ $=$ $p_{0}+\epsilon p_{1},$ (5)
$\Psi$ $=$ $\Psi_{0}+\epsilon\Psi_{1},$ (6)
we obtain the wave equation
$-\partial_{t}\biggl{[}\frac{\partial\rho}{\partial p}\rho_{0}(\partial_{t}\Psi_{1}+\mathbf{v}_{0}\cdot\nabla\Psi_{1})\biggr{]}$
$+\nabla\cdot\biggl{[}\rho_{0}\nabla\Psi_{1}-\frac{\partial\rho}{\partial p}\rho_{0}\mathbf{v}_{0}(\partial_{t}\Psi_{1}+\mathbf{v}_{0}\cdot\nabla\Psi_{1})\biggr{]}=0,$ (7)
where the local speed of sound, $c_{s}$, is defined by
$c_{s}^{-2}\equiv\frac{\partial\rho}{\partial p}.$ (8)
The wave equation (7) describes the propagation of the linearized scalar potential $\Psi_{1}$; that is, it governs the propagation of the phase fluctuations as weak excitations in a fluid’s surface that can be rewritten as a curved-spacetime wave equation
$\frac{1}{\sqrt{-g}}\partial_{\mu}(g^{\mu\nu}\sqrt{-g}\partial_{\nu}\Psi_{1})=0.$ (9)
where the effective curved spacetime $g_{\mu\nu}$ has a line element of the form
$ds^{2}=\frac{\rho_{0}}{c_{s}}\left[-c_{s}^{2}dt^{2}+(dx^{i}-v_{0}^{i}dt)\delta_{ij}(dx^{j}-v_{0}^{j}dt)\right].$ (10)
Unruh assumed the background flow as a spherically-symmetric, stationary inviscid fluid. By including circulation to the fluid flow, the acoustic metric given by Eq. (10) obtains an explicit form in cylindrical coordinates, such as
$ds^{2}=$ $\frac{\rho_{0}}{c_{s}}\biggl{\{}-[c_{s}^{2}-(v_{0}^{r})^{2}-(v_{0}^{\theta})^{2}]dt^{2}+\frac{c_{s}^{2}}{c_{s}^{2}-(v_{0}^{r})^{2}}dr^{2}$
$+r^{2}d\theta^{2}+dz^{2}-2v_{0}^{\theta}rdtd\theta\biggr{\}},$ (11)
where the following coordinate transformations were performed
$dt\rightarrow dt-\frac{v_{0}^{r}}{c_{s}^{2}-(v_{0}^{r})^{2}}dr,$ (12)
$d\theta\rightarrow d\theta-\frac{v_{0}^{\theta}v_{0}^{r}}{r[c_{s}^{2}-(v_{0}^{r})^{2}]}dr,$ (13)
with $\vec{v}=v_{0}^{r}\hat{r}+v_{0}^{\theta}\hat{\theta}$. Since we included rotation, the equation $c_{s}^{2}-(v_{0}^{r})^{2}-(v_{0}^{\theta})^{2}=0$ provides a condition for finding the effective ergosphere radius, which in turn are connected with energy extraction from BHs, and superradiance [70]. The radial and angular components of the fluid flow velocity can be written as
$v_{0}^{r}=-\frac{D}{r},$ (14)
$v_{0}^{\theta}=\frac{C}{r},$ (15)
where $D$ and $C$ can be identified as the radial draining inflow parameter and the circulation, respectively. Therefore, we can write the line element for rotating acoustic BHs as
$ds^{2}=$ $-\biggl{[}f(r)-\frac{C^{2}}{r^{2}}\biggr{]}dt^{2}+\frac{1}{f(r)}dr^{2}$
$+r^{2}d\theta^{2}+dz^{2}-2Cdtd\theta,$ (16)
where the acoustic lapse function $f(r)$ is given by
$f(r)=1-\frac{D^{2}}{r^{2}}.$ (17)
Note that we have dropped a position-independent prefactor by fixing $c_{s}=1$ and assumed a constant fluid density $\rho_{0}$. The ergosphere of this background can be found by solving
$1-\frac{D^{2}}{r^{2}}=\frac{C^{2}}{r^{2}},$ (18)
which has a unique real solution given by
$r_{\rm erg}=\sqrt{D^{2}+C^{2}}.$ (19)
The acoustic event horizons form once the radial component of the fluid flow velocity exceeds the speed of sound, that is, when
$f(r)=0,$ (20)
which has two real solution given by
$r_{\pm}=\pm D.$ (21)
Thus, $r=r_{+}$ designates a future acoustic event horizon (acoustic BH), while $r=r_{-}$ depicts a past acoustic event horizon (acoustic white hole). In addition, the dragging angular velocity of the acoustic horizon, $\Omega$, is given by
$\Omega(r)\equiv-\frac{g_{03}}{g_{33}}=\frac{C}{r^{2}},$ (22)
such that $\Omega_{+}=\Omega(r_{+})=C/r_{+}^{2}$ is the frame dragging at the acoustic event horizon $r=r_{+}$.
III Horizonless vortex fluid flow
When one nullifies the drain, i.e. when $D=0$, then the vortex becomes purely rotational and radial fluid flow is absent. In this case, the background presents no acoustic event horizon, no radial inflow, and hence it describes a horizonless vortex fluid flow; an effective geometry that has been recently constructed in the laboratory with the use of superfluid Helium [68].
A $(2+1)$-dimensional metric describing such a vortex model can be written as
$ds^{2}=-\biggl{(}1-\frac{C^{2}}{r^{2}}\biggr{)}dt^{2}+dr^{2}+r^{2}d\theta^{2}-2Cdtd\theta,$ (23)
where we have set $z=h_{0}=\mbox{constant}$. In what follows, we will analyze the propagation of scalar excitations at the region of the general vortex spacetime given by Eq. (23), with $D=0$, and calculate the bound spectra. We will adopt an approach that includes the transformation of the equation of motion onto a Mathieu differential equation. We will then use the possible pairs of boundary conditions to obtain bound states (BSs) numerically, compare them with respect to different boundary conditions, and finally try to comprehend if there is a boundary condition that is more physical and closer to the giant quantum vortex experiment [68].
III.1 Analogy between the effective metric used and modern laboratory experiments: an interlude
The metric depicted by Eq. (16) describes a rotating acoustic BH, with an acoustic horizon at $r=r_{+}=D$, where the rotational velocity of the fluid is imprinted in the circulation parameter $C$. The way that Eq. (16) is set-up allows for two tunable configurations; the typical rotating acoustic BH ($D>0$) or a rotating horizonless vortex without an acoustic horizon ($D=0$), i.e., a vortex without any radial inflow of surface fluctuations towards its center. The latter case presents a perfect testbed for contemporary laboratory experiments with fluids, superfluids and lasers.
In particular, the experimental setup that we will try to qualitatively replicate from Refs. [68, 22] has a lot of similarities with the geometry described by Eq. (23). There, they investigate phenomena in the limit of negligible viscosity by using cooled-down superfluid 4He. Specifically, at $T=1.75$ K, at which the experiment takes place, the kinematic viscosity of the superfluid is reduced by a factor of $100$ with respect to water experiments. The key feature of superfluid 4He is the fact that it supports topological defects known as quantum vortices that carry circulation quanta. Each circulation quantum forms an irrotational (zero-curl) flow field in its vicinity. Due to such discretization, a vortex of superfluid 4He can manifest itself only as a giant (multiply quantized) vortex. Such vortices can mimic rigid-body rotation and altogether suggest that an extensive draining vortex of 4He is a feasible candidate for experimental simulations of a quantum-field theory in curved spacetime. An important component of the setup in [68, 22] is a spinning propeller, which establishes a continuous circulation. We can loosely identify the propeller frequency with the parameter $C$ in Eq. (23). Even though the resulting circulation stems from circulation quanta, a record-breaking amount of these quanta have been used ($\sim 4\times 10^{4}$) that allows for the quantization of circulation to be disregarded; the giant quantum vortex is effectively classical. The most significant aspect of the apparatus in [68, 22] is the absence of radial flow of interface fluctuations. This means that the experiment includes pure circulation and no acoustic horizon, as it is the case for the metric in Eq. (23), when $D=0$.
Note that the geometry used here assumes the hydrodynamic limit, which effectively cancels any potential nonlinearities that emerge from dispersion. Not taking into consideration dispersive, nonhydrodynamic, effects practically leads to Unruh’s historical acoustic metric proposal [11] that formed the onset of the interest in hydrodynamic analog models of gravity and the construction of laboratory experiments. Nevertheless, avoiding dispersion in laboratory experiments is extremely difficult since nonlinearities are always part of realistic analog experiments [33, 71, 47, 72, 73, 68, 22]. Even so, simplified BH analog metrics like the one used here, can competently simulate the basic features and dynamics that take place in realistic analog BH experiments.
In this work, we will qualitatively replicate some results that have been obtained experimentally in [68], regarding interface fluctuations in circulating superfluid Helium and the resulting QBSs. We note that our perturbative scheme involves acoustic, non-dispersive, waves which is quite different from the dispersive interface waves of [68]. This is due to the use of a superfluid. Our metric is independent of the fluid’s state or in general the state of matter used to create the vortex. Therefore, it can be applied for other experiments as well, though here we will commence the study by testing the robustness of our analysis by comparing our results with those found in [68].
IV The bound spectrum of a horizonless rotating vortex
IV.1 Scalar wave equation
We are interested in the basic features of the effective geometry given by Eq. (23), in particular the propagation of excitations on the surface of the effective fluid flow. In order to perform this analysis, we have to solve the wave equation (9) with appropriate boundary conditions to obtain the QBSs. By substituting Eqs. (23) into Eq. (9), we obtain the following scalar wave equation
$-\frac{\partial^{2}\Psi_{1}}{\partial t^{2}}+\frac{1}{r}\frac{\partial}{\partial r}\biggl{(}r\frac{\partial\Psi_{1}}{\partial r}\biggr{)}+\biggl{(}\frac{1}{r^{2}}-\frac{C^{2}}{r^{4}}\biggr{)}\frac{\partial^{2}\Psi_{1}}{\partial\theta^{2}}-\frac{2C}{r^{2}}\frac{\partial^{2}\Psi_{1}}{\partial t\partial\theta}=0,$ (24)
where the scalar wave function is such that $\Psi_{1}=\Psi_{1}(t,r,\theta)$. Next, since the line element given by Eq. (23) resembles those of rotating spacetimes, we can use the following separation ansatz for the perturbation $\Psi_{1}$
$\Psi_{1}\simeq\mbox{e}^{-i\omega t}R(r)\mbox{e}^{im\theta},$ (25)
where $\omega$ is the frequency (energy, in the natural units), $R(r)$ is the radial function, and $m$ is the azimuthal number (a discrete integer running from $\infty$- ∞ to +$\infty$+ ∞). By substituting Eqs. (23)-(25) into Eq. (9), we obtain a radial master equation given by
$R^{\prime\prime}(r)+\frac{1}{r}R^{\prime}(r)+\biggl{(}\omega^{2}-\frac{m^{2}+2Cm\omega}{r^{2}}+\frac{C^{2}m^{2}}{r^{4}}\biggr{)}R(r)=0.$ (26)
IV.2 Bound states
Equation (26) can be re-casted into the form of a Mathieu differential equation [74] in order to be solved for its eigenvalues and eigenfunctions. We define a new radial coordinate $x$ such that
$r=\tau\mbox{e}^{ix}\,,\quad\quad x\in\mathbb{R}.$ (27)
By substituting (27) into (26) we obtain
$R^{\prime\prime}(x)+\biggl{(}m^{2}+2Cm\omega-\frac{C^{2}m^{2}\mbox{e}^{-2ix}}{\tau^{2}}-\tau^{2}\omega^{2}\mbox{e}^{-2ix}\biggr{)}R(x)=0,$ (28)
or, after some algebraic calculations
$R^{\prime\prime}(x)+[m(m+2C\omega-2C\omega\cos(2x)]R(x)=0,$ (29)
where we have defined
$\tau=\sqrt{\frac{Cm}{\omega}}.$ (30)
This transformation can be understood as a mapping from the real variable $x\in\mathbb{R}$ to a complex-valued radial coordinate $r(x)$ lying on a circle of radius $\tau$ in the complex plane. Thus, this transforms Eq. (26) to the following normal form
$R^{\prime\prime}(x)+[a-2q\cos(2x)]R(x)=0.$ (31)
That is the Mathieu equation [75], whose general solution is given by
$R(x)=C_{1}\ \mbox{MathieuC}(a,q,x)+C_{2}\ \mbox{MathieuS}(a,q,x),$ (32)
where $C_{1}$ and $C_{2}$ are constants to be determined, MathieuC and MathieuS are the even and odd Mathieu functions, respectively, and the parameters $a$ and $q$ are given by
$a$ $=$ $m^{2}+2Cm\omega,$ (33)
$q$ $=$ $Cm\omega.$ (34)
To obtain the eigenfrequencies, we need to impose appropriate boundary conditions on the radial solution given by Eq. (32). The boundary value problem related to Mathieu functions will define the physical meaning of the solutions with respect to an “idealized” version of the laboratory experiment we are trying to qualitatively reproduce. In fact, since the Mathieu functions are periodic, the solutions we will obtain from our idealized experiment will be BSs, instead of QBSs. Nevertheless, we will see that to a certain extent, these solutions describe the qualitative aspects of the dynamics of the actual laboratory experiment in [68].
Figure 1: Two-dimensional wave spectra for co-rotating and counter-rotating BSs for varying discrete azimuthal number $m$. We show the first three (non-zero) modes with unitary circulation $C=1$. Even though the subfigures depict continuous curves for the BSs, these are just the discrete BS frequencies for different $m$, that we joined together in a fit.
The Mathieu functions become periodic for some characteristic values of $a$ for given $q$. In any half-open interval of length $\pi$ on the real axis, an even Mathieu function with $n$ zeros in the interval $0\leq x<\pi$, $\mbox{MathieuC}_{n}(q,x)$, has characteristic values denoted by $A_{n}(q)$, with $n=0,1,2,\ldots$, while an odd Mathieu function, $\mbox{MathieuS}_{n}(q,x)$, has characteristic values denoted by $B_{n}(q)$, with $n=1,2,\ldots$. These characteristic values satisfy any of the following pairs of boundary conditions [75]:
(I) $\qquad R(0)=R(\pi)=0,\ \mbox{for}\ \mbox{MathieuS}_{n}(q,x)\ \mbox{with period of}\ \pi;$ (35)
(II) $\qquad R^{\prime}(0)=R^{\prime}(\pi)=0,\ \mbox{for}\ \mbox{MathieuC}_{n}(q,x)\ \mbox{with period of}\ \pi.$ (36)
On the other hand, if $R(x)$ is either $\mbox{MathieuS}_{n}(q,x)$ or $\mbox{MathieuC}_{n}(q,x)$, then $R(x)$ and $R(\pi-x)$ satisfy the same differential equation and the same boundary conditions, and must be constant multiples of each other so that $R(x)$ is either an even or an odd function of $\pi/2-x$. This leads to four additional pairs of boundary conditions [75]:
(III) $\qquad R(0)=R(\pi/2)=0,\ \mbox{for}\ \mbox{MathieuS}_{2n+2}(q,x)\ \mbox{with period of}\ \pi;$ (37)
(IV) $\qquad R(0)=R^{\prime}(\pi/2)=0,\ \mbox{for}\ \mbox{MathieuS}_{2n+1}(q,x)\ \mbox{with period of}\ 2\pi;$ (38)
(V) $\qquad R^{\prime}(0)=R(\pi/2)=0,\ \mbox{for}\ \mbox{MathieuC}_{2n+1}(q,x)\ \mbox{with period of}\ 2\pi;$ (39)
(VI) $\qquad R^{\prime}(0)=R^{\prime}(\pi/2)=0,\ \mbox{for}\ \mbox{MathieuC}_{2n}(q,x)\ \mbox{with period of}\ \pi.$ (40)
Here, for each $n=0,1,2,\ldots$ there is exactly one characteristic function of each of these four boundary value problems, and $n$ is the number of zeros in the interval $20<x<\pi/$ 2.
Figure 2: Amplitudes of acoustic waves $|R^{(i)}_{mn}(r)|$ corresponding to the $m=1$ mode, with unitary circulation $C=1$. Here, we plot the first three (non-zero) overtones of the BSs, with respect to the radius $r$ ranging from 0 to $\pi$ and frequencies $\omega^{(i)}_{mn}$, for all possible boundary conditions (i=I-VI). We also show the effective potential $V^{(i)}_{{\rm eff},mn}$ with gray curves [the $\omega$-dependent factor, multiplying $R(r)$ in Eq. (26)] for different BS frequency overtones. The solid colored lines correspond to the amplitude of rescaled BSs, which show a characteristic oscillatory pattern, that can be identified with their corresponding BS frequencies (dashed colored lines) outside the potential. The leftmost vertical dashed line corresponds to a reference point at $r=0.8$, while the rightmost one corresponds to $r=\pi$, as seen in the left subfigure of the effective potential, and $r=0.8$ as seen in the right subfigure of the BS amplitudes. Crossing points with the effective potential are marked by gray points.
Figure 3: Comparison between the absolute value of the wave eigenfunctions $|R^{(i)}_{mn}(x)|$, in the $x$-space of Mathieu functions, for all possible boundary conditions (i=I-VI). Here, the azimuthal number is $m=1$ and the circulation is $C=1$, where we vary the overtone number $n$.
From a purely mathematical perspective, we cannot exclude any of the boundary conditions. Thus, on the solution for the radial part of the massless Klein-Gordon equation in a vortex rotational fluid flow, given by Eq. (32), we will impose all six pairs of boundary conditions, so that the parameter $a$ becomes characteristic value of one of the periodic Mathieu functions, i.e.,
$a=B_{N}(q),\ \mbox{for}\ \mbox{MathieuS}_{N}(q,x),$ (41)
or
$a=A_{N}(q),\ \mbox{for}\ \mbox{MathieuC}_{N}(q,x).$ (42)
[…]
Figure 1 presents the theoretical predictions of the minimum frequency permissible for propagation of BSs for a given circulation $C$, and a variety of azimuthal numbers $m$ and overtones $n$. We consider both co-rotating ($m>0$) and counter-rotating ($m<0$) BS solutions, from which the symmetry $\omega_{mn}\rightarrow-\omega_{-mn}$ arises. We note that this symmetry holds in our analysis due to the idealization of the experiment, that lacks nonlinear dispersion relations. This symmetry holds for all overtone number $n$. Notice that in some cases of boundary conditions, the discrete BS frequency changes from positive to negative, though since they are purely real, the periodicity of the solution is still preserved. By comparing our resulting spectra with the experimental observations in [68] (see Figs. 2f and 2g), we find that their behavior for different boundary conditions remains qualitatively the same with those found in [68], especially for radii of the superfluid interface that are much larger than the experimental potential barrier maximum. In fact the larger the radius, the more the aforementioned symmetry approximately holds in this particular experiment due to the reduction of the angular velocity distribution. Of course, in our case the choice of $C$ points to a velocity distribution, therefore when choosing $C$, we effectively change the radius of the interface, while ignoring other effects, such as dissipation of the wave’s energy by viscous damping and by scattering into the vortex. Nevertheless, this is somewhat balanced by the stochastic drive originating from the fluid flow and mechanical vibrations in the experimental setup [68], thus our assumptions and resulting spectral space $(m,f)$ of discrete BSs are not necessarily purely theoretical. Another similarity that appears between our data in Fig. 1 and those in Fig. 2 of [68] is the fact that only a certain region of the spectral space is populated with BSs. We, also, observe that only some high-frequency (equivalent to high-energy) waves have the capability to propagate.
Figure 2 shows the BS eigenfunctions with respect to the radial coordinate $r$. The demonstration unveils the pattern of the BSs with respect to the radius of the idealized vortex fluid flow, for a fixed circulation $C=1$. The shapes of the effective potentials, for different overtone numbers $n$, are more pronounced closer to their respective peaks. Different boundary conditions give rise to different BS eigenfunctions, and respective BS frequencies, as well as higher- or lower-frequency maxima of the effective potential peaks. The most pronounced peak appears for the boundary condition (V).
Interestingly, similar co-rotating BSs (dubbed QBSs) have been found in the laboratory experiment involving a giant quantum vortex of superfluid 4He [68]. Even though there is a resemblance between our results and those extracted from the realistic experiment, this is only a qualitative resemblance. Our effective spacetime metric does capture the important qualities and features of the realistic experiment, but it is not yet suitable to draw solid quantitative conclusions, especially because the patterns that appear in actual superfluid vortex experiments, as in [68], regard interface waves, and not acoustic waves as the idealized manifestation of the effective metric and analysis involves. Our results, therefore, are qualitatively robust and may be of significance in the future when all aspects of the experimental setup are taken into account, such as nonlinear dispersion, interface waves, the transition between solid and hollow core vortices and, finally, which choice of boundary conditions is the most suitable one for such an experiment. Our analysis has also been performed with respect to the Mathieu-space radial coordinate $x$, in order to better understand if the boundary conditions of the BS eigenfunctions are met.
Figure 3 demonstrates the absolute values of BS eigenfunctions for the six pairs of boundary conditions with respect to the coordinate $x$. We focus on the co-rotating $m=1$ mode, with circulation set to $C=1$, though these choices are practically random since qualitatively similar results appear for higher $m$ and varying $C$. We observe that in the $x$-space, all the boundary conditions, that are imposed in order to obtain the BS solutions, are appropriately satisfied.
The most important question then is which boundary condition is more physical for the calculation of BSs in their domain of existence. To find QBSs of BHs, the boundary conditions imposed are purely ingoing waves at the BH event horizon and purely reflecting waves at infinity [76]. Here, the experiment we are trying to reproduce possesses no radial inflow of perturbations into the vortex, therefore there is no acoustic event horizon such as those appearing in BHs and BH analogs. Thus, a different boundary condition has to be imposed at the “center” of the vortex $x=0$. Since there is no event horizon, there is also an absence of a singularity at the internal vicinity of the vortex. This leads to the typical boundary condition that is imposed close to the center of stars when their perturbations are studied, i.e. regularity at the central region of the star [77, 78, 79, 80, 81, 82]. The boundary conditions at the center ensure that solutions remain finite and physically plausible. This is also the mathematical requirement to avoid singularities.
One of the most general case of boundary conditions that ensures regularity is the Neumann boundary condition that specifies the values of the derivative applied at the boundary of a domain, which in our case translates to $R^{\prime}(x=0)=\textrm{const.}$. in order for the solutions to remain finite there. This means that there is no flow across the boundary, effectively making the central region a no-flux zone for the perturbation. In this context, it essentially implies that the perturbation is symmetric, as it occurs with solutions to the Mathieu equation which have certain periodicity. The normal derivative of the perturbed quantity (say the temperature or the density of the superfluid in [68]) is set to zero or a constant, meaning there is no change in the slope of the quantity across the internal region.
There are more restrictive boundary conditions, such as the Dirichlet one where the field satisfies $R(x=0)=\textrm{const.}$. at the central region, that is there is no perturbation at the vicinity of the core. Nevertheless, the constant resulting from the Neumann boundary condition can be chosen to be zero or otherwise, in order to cover wider scenarios of regularity, such as regularity at the inner (close to the central) region’s boundary and the outer boundary of the rotational fluid that is deemed to possess only an azimuthal velocity distribution. From a mathematical perspective, the boundary conditions that we can apply in Mathieu differential equations either set the field or its normal derivative equal to zero at $x=0$ [see Eqs. (35)-(40)]. In principle, both a Dirichlet or a Neumann boundary condition can be used here, since they are both physically viable. Nevertheless, we could argue that in order to keep the temperature or the density of the superfluid constant at the central region of the vortex (and not zero, according to [68]), then the most general boundary condition is the Neumann one that ensures both regularity and a non-zero finite constant around $x=0$. From our perspective, this narrows down the possible boundary conditions of the Mathieu equation to (36), (39) and (40).
As mentioned above, to find QBSs (or BSs) in BHs (or horizonless compact objects) we impose purely reflecting boundary conditions at infinity. This corresponds to Dirchlet boundary conditions at infinity, and in our case the eigenfunction should become zero at the boundary of the Mathieu space that our solution exists. Following our argumentation, the pair of boundary conditions needed in order for the resulting BSs to have a more appropriate physical meaning are the ones described by (39). This coincides with the boundary condition (V) that allows for the highest frequency peak of the effective potential, which coincidentally looks more similar to what was found in the experimental apparatus of Ref. [68]. This, of course, is not a statement of quantitative agreement between the metric’s perturbative behavior and the experimental data, but rather a mere coincidence. In principle, all boundary conditions of the Mathieu equation can qualitatively reproduce the shapes and overall structure shown in Figs. 2, 4, and 5 of [68].
V Final remarks
We have proposed the use of an acoustic metric that describes a draining and rotational fluid flow, specified by the drain $D$ and the circulation $C$ parameters included in the metric, in order to qualitatively reproduce current experimental data. This acoustic spacetime has been constructed assuming the hydrodynamic limit, that discards any nonlinear dispersion relations between the velocity density profile and the perturbations that we study. There are an abundance of applications of this metric in order to mimic analog BH experiments. A special case of interest in this work is the giant quantum vortex, recently constructed in the laboratory [68]. This vortex has two configurations; solid or hollow core. Even though the utilized metric does not directly discusses the possibility of two different cores, the BS analysis of acoustic waves shows a lot of qualitative similarities with the actual experiment and seems universal for BSs, irregardless of the core structure, since there is no radial flow of perturbations in the experiment and in the effective metric ($D=0$). By obtaining a wave equation for the propagation of acoustic waves in the analog spacetime, we find a transformation that turns it into a Mathieu differential equation. Therefore, we find the BS solutions and frequencies through the Mathieu functions by applying all possible boundary conditions that these functions must, mathematically, satisfy. We then discuss their structure and compare them with the quasi-periodic interface BSs found in the vortex experiment. Our findings indicate that the phenomenology of such an intricate experiment could be qualitatively reproduced by our extremely simplified metric and perturbation analysis.
We believe that our results for the BSs of acoustic waves might be helpful to the experimental community, since they can describe periodic acoustic waves propagating in a rotational fluid flow without radial inflow. In this sense, we expect that one could test the spacetime background under consideration and hence determine, in principle, if and how the idealized acoustic metric used here can provide more than just qualitative data.
To that end, we need to consider a key ingredient that is not directly discussed in the horizonless tunable acoustic rotational analog. This is the ability to also tune the horizonless metric between a solid and a hollow core vortex. In order to be able to perform such a change to the metric, we need to include a new parameter and/or a functional of an existing parameter that enables us to perform this transition. The effective analog metric includes, besides the drain parameter $D$, a parameter for the circulation $C$ of the vortex. $C$ in turn, traces back to an angular velocity profile and a corresponding propeller frequency that controls the absence or presence of a hollow core. Thus, it might be the case that the circulation can be included in the metric in a form of a functional that enables this transition, depending on its value. This work needs to take advantage of the experimental data from Ref. [68], in order to build (fit) a relation between various $C$ and the corresponding propeller frequencies of the experimental apparatus. We leave this work for the future.
Another interesting feature that can emerge from such experiments, and the resulting idealized metric under consideration, is the fact that there is no inflow of radial perturbations into the core. In a sense, the experiment may serve as a proxy for astrophysical environments; either extremely massive ones that surround galaxies, where the inflow of matter is negligible with respect to the lifetime of the environment (such as dark matter halos) or with dispersion relations that describe high-energy trans-Planckian effects around BHs. Hence, our analysis, and the successful control of the experiment in [68], may lead the way for a more simplistic, yet accurate enough, design of decoupling environmental effects (matter perturbations) from gravitational ones (especially polar gravitational perturbations in astrophysical BHs), and in general the relativistic treatment of astrophysical environments around BHs [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 90, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109].
Acknowledgements.
The authors are indebted to Prof. Kostas D. Kokkotas for fruitful discussions and comments in the manuscript. The authors also would like to thank Pietro Smaniotto and Leonardo Solidoro for their help in better understanding the experimental apparatus and its function. This study was financed in part by the Conselho Nacional de Desenvolvimento Científico e Tecnológico – Brasil (CNPq) – Research Project No. 440846/2023-4 and Research Fellowship No. 201221/2024-1. H.S.V. is partially supported by the Alexander von Humboldt-Stiftung/Foundation (Grant No. 1209836). Funded by the Federal Ministry of Education and Research (BMBF) and the Baden-Württemberg Ministry of Science as part of the Excellence Strategy of the German Federal and State Governments – Reference No. 1.-31.3.2/0086017037. K.D. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., under the Scientific Employment Stimulus – Individual Call – Grant No. 2023.07417.CEECIND/CP2830/CT0008. This project has received funding from the European Union’s Horizon-MSCA-2022 research and innovation programme “Einstein Waves” under grant agreement No. 101131233.